# Emphasizing Discriminative Features for Dataset Distillation in Complex Scenarios

### v1

@@ C.1 Cross-architecture Evaluation
C.2 Eval. without Knowledge Distillation
Starting from Wang et al. (2020), representative dataset distillation (DD) methods (Zhao et al., 2021; Zhao & Bilen, 2021b; Cazenavette et al., 2022; Wang et al., 2022) establish a general workflow as follows: 1) Distillation: At this stage, information from the real dataset is fully accessible to the DD algorithm to train synthetic data. 2) Evaluation: After the distilled dataset is obtained, the evaluation is performed by training a randomly initialized model on the distilled data. Specifically, in the context of classification, the objective is to minimize cross-entropy loss. Recently, some new methods (Yin et al., 2023; Sun et al., 2023) introduced teacher knowledge into the student model by applying knowledge distillation. Although it helps improve performances to a large extent, it may not be able to reflect the effectiveness of dataset distillation accurately.
@@
C.3 Distorted Images of Large Enhancement Factor
@@
C.4 Changing Trends of Discriminative Areas
@@
Appendix D Comp-DD Benchmark
D.1 Subset Details
@@ D.2 Complexity Metrics
D.3 Benchmark Hyper-parameters
@@
Hyper-parameters $\alpha$ $\beta$ $K$ $T$ batch_syn lr_pixel lr_label syn_steps
CDD-Bird-Easy 1 0 1 50 10000 1000 10000 2.0 40
10 0.25 1 100 400 1000 3.0
50 0.375 2 200 200 100 5.0
CDD-Bird-Hard 1 0 1 50 10000 1000 10000 2.0 40
10 0.25 1 100 400 1000 3.0
50 0.375 2 200 200 100 5.0
CDD-Dog-Easy 1 0 1 50 10000 1000 10000 2.0 40
10 0.25 1 100 400 1000 5.0
50 0.375 2 200 200 100 5.0
CDD-Dog-Hard 1 0 1 50 10000 1000 10000 2.0 40
10 0.25 1 100 400 1000 2.0
50 0.375 2 200 200 100 5.0
CDD-Car-Easy 1 0 1 50 10000 1000 10000 3.0 40
10 0.25 1 100 400 1000 3.0
50 0.375 2 200 200 100 5.0
CDD-Car-Hard 1 0 1 50 10000 1000 10000 3.0 40
10 0.25 1 100 400 1000 3.0
50 0.375 2 200 200 100 5.0
@@
Appendix E Visualization of Distilled Images on ImageNet-1K
@@
Appendix F More Related Work
In Table 13, we present a comprehensive summary of previous dataset distillation methods, categorized by different approaches. There are four main categories of dataset distillation: gradient matching, trajectory matching, distribution matching, and generative model-based methods. Recently, some works (Yin et al., 2023; Sun et al., 2023; Yu et al., 2024) add knowledge distillation during the evaluation stage of dataset distillation.
@@ Appendix F More Related Work
Kernel-based KIP-FC (Nguyen et al., 2021a)
KIP-ConvNet (Nguyen et al., 2021b)
FRePo (Zhou et al., 2022)
RFAD (Loo et al., 2022)
RCIG (Loo et al., 2023)
Gradient-matching DC (Zhao et al., 2021)
DSA (Zhao & Bilen, 2021b)
DCC (Lee et al., 2022a)
LCMat (Shin et al., 2023)
Trajectory-matching MTT (Cazenavette et al., 2022)
Tesla (Cui et al., 2022a)
FTD (Du et al., 2022)
SeqMatch (Du et al., 2023)
DATM (Guo et al., 2024)
ATT (Liu et al., 2024)
NSD (Yang et al., 2024)
PAD (Li et al., 2024)
SelMatch (Lee & Chung, 2024)
Distribution-matching DM (Zhao & Bilen, 2021a)
CAFE (Wang et al., 2022)
IDM (Zhao et al., 2023)
DREAM (Liu et al., 2023b)
M3D (Zhang et al., 2023)
Generative model DiM Wang et al. (2023)
GLaD (Cazenavette et al., 2023)
H-GLaD (Zhong et al., 2024)
LD3M (Moser et al., 2024)
IT-GAN (Zhao & Bilen, 2022)
D4M Su et al. (2024)
Minimax Diffusion Gu et al. (2023)
+ Knowledge distillation for evaluation SRe2L (Yin et al., 2023)
RDED (Sun et al., 2023)
HeLIO (Yu et al., 2024)

### v2

@@ 11.1 Cross-architecture Evaluation
11.2 Eval. without Knowledge Distillation
Starting from [54], representative dataset distillation (DD) methods [71, 69, 1, 51] establish a general workflow as follows: 1) Distillation: At this stage, information from the real dataset is fully accessible to the DD algorithm to train synthetic data. 2) Evaluation: After the distilled dataset is obtained, the evaluation is performed by training a randomly initialized model on the distilled data. Specifically, in the context of classification, the objective is to minimize cross-entropy loss. Recently, some new methods [64, 48, 74, 73] introduced teacher knowledge into the student model by applying knowledge distillation. Although it helps improve performances to a large extent, it may not be able to reflect the effectiveness of dataset distillation accurately.
@@
IPC Grad-CAM Model
ConvNetD5 ResNet18 ResNet50 VGG11
1 52.3 52.6 52.5 52.5
10 71.2 71.0 70.8 70.7
50 77.4 77.8 77.6 77.6
Table 13: Results of using different Grad-CAM models on ImageNette. The choice of model only has minor influence on the performance.
IPC 1 10 50 200 300
Latency (s) 0.63 0.52 0.74 0.68 0.94
Table 14: Latency of extracting Grad-CAM activation maps using ResNet18. For each IPC in our experiments, the latency is less than one second.
11.3 Distorted Images of Large Enhancement Factor
@@
11.4 Changing Trends of Discriminative Areas
@@
11.5 Impact of different Grad-CAM Models
In our experiments, we use ResNe18 as the Grad-CAM model to extract activation maps. However, the choice of Grad-CAM model does not have a significant impact on the performance, as long as it has been trained on the target dataset. As shown in Table 13, differences between performances of different Grad-CAM models are within 0.5, demonstrating that our discriminative area enhancement module doesn’t depend on the choice of Grad-CAM model.
11.6 Latency of Grad-CAM
We use torchcam [11] implementation of various Grad-CAM methods to extract activation maps. In Table 14, we show the latencies of extracting activation maps of various IPCs. Notably, these latencies are all below one second, demonstrating a high inference speed. In our experiments, the maximum number of extractions for one distillation is 200 (on IPC1). Thus, the total time used for activation map extraction is at most two minutes, which is neglectable compared with the latency of the full distillation (several hours). In conclusion, our use of Grad-CAM activation maps to provide guidance doesn’t reduce the efficiency of the backbone.
@@ 11.6 Latency of Grad-CAM
12 Comp-DD Benchmark
12.1 Subset Details
@@ 12.2 Complexity Metrics
12.3 Benchmark Hyper-parameters
@@
Hyper-parameters $\alpha$ $\beta$ $K$ $T$ batch_syn lr_pixel syn_steps
CDD-Bird-Easy 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 400 1000
50 0.375 2 200 200 100
CDD-Bird-Hard 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 400 1000
50 0.375 2 200 200 100
CDD-Dog-Easy 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 400 1000
50 0.375 2 200 200 100
CDD-Dog-Hard 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 400 1000
50 0.375 2 200 200 100
CDD-Car-Easy 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 400 1000
50 0.375 2 200 200 100
CDD-Car-Hard 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 400 1000
50 0.375 2 200 200 100
@@
13 Visualization of Distilled Images on ImageNet-1K
@@
14 More Related Work
In Table 16, we present a comprehensive summary of previous dataset distillation methods, categorized by different approaches. There are four main categories of dataset distillation: gradient matching, trajectory matching, distribution matching, and generative model-based methods. Recently, some works [64, 48, 65] add knowledge distillation during the evaluation stage of dataset distillation.  [40, 33, 4, 36, 63, 29, 66, 10, 58, 55, 50, 17, 35, 34, 62, 7, 57, 56, 60, 5].
@@ 14 More Related Work
Kernel-based KIP-FC [38]
KIP-ConvNet [39]
FRePo [78]
RFAD [31]
RCIG [32]
Gradient-matching DC [71]
DSA [69]
DCC [22]
LCMat [44]
Trajectory-matching MTT [1]
Tesla [6]
FTD [8]
SeqMatch [9]
DATM [13]
ATT [27]
NSD [61]
PAD [25]
SelMatch [24]
Distribution-matching DM [68]
CAFE [51]
IDM [72]
DREAM [30]
M3D [67]
NCFD [53]
Generative model DiM [52]
GLaD [2]
H-GLaD [77]
LD3M [37]
IT-GAN [70]
D4M [46]
Minimax Diffusion [12]
+ Knowledge distillation for evaluation SRe2L [64]
RDED [48]
HeLIO [65]
Others MIM4DD [43]
DQAS [75]
LDD [76]